When invested at an annual interest rate of 7% the amount earned P11,200 of simple interest in 2 years. How much money was originally invested?

1
Expert's answer
2022-01-17T17:57:22-0500

for simple interest:

"A=P(1+rt)"


 originally invested:

"P=A/(1+rt)=11200/(1+0.07\cdot2)=9824.56"